OpenStudy (anonymous):

Write an expression for a) an unknown number subtracted by 17 ( just the answer) b) perimeter of a rectangle with base 3x + 9 and the height one-third of the base length. Simplify even the expression! Present your solution.

OpenStudy (nirmalnema):

17-x

OpenStudy (nirmalnema):

2(h+b) b=3x+9 h=(3x+9)/3 =3(x+3)/3 =x+3 so 2(3x+9+x+3)=(4x+12)2 =8x+24

OpenStudy (anonymous):

so do i write the answers like this? A) 17-x?? and B) 2(h+b) b=3x+9 h=(3x+9)/3 =3(x+3)/3 =x+3 so 2(3x+9+x+3)=(4x+12)2 =8x+24

OpenStudy (nirmalnema):

yup..

OpenStudy (anonymous):

in b) , where does the number 2 comes from? ... omg, i suck..

OpenStudy (nirmalnema):

its the formula.. 2(h+b)

OpenStudy (nirmalnema):

perimeter of recangle

OpenStudy (anonymous):

O.o i suck ! lol thank u as always nirmalnema!!! <3

OpenStudy (nirmalnema):

welcome